### v2

# Ultrafast Electron Temperature Dynamics in Spintronic Terahertz Emitters Studied by Optical-Pump Terahertz-Probe spectroscopy

Abstract
Spintronic terahertz emitters pumped by femtosecond lasers offer an increasingly popular approach to generate broadband terahertz radiation. The strength of the emitted terahertz field is limited in part by the optical damage threshold at the pump wavelength of the spintronic terahertz emitter. Their thermal management can be improved by better understanding the electron temperature relaxation in the spintronic metal layer. Here we present a measurement of the electron temperature dynamics on a picosecond timescale using optical-pump terahertz-probe spectroscopy. We observe that the optical pump induces a change in the terahertz transmissivity of the spintronic terahertz emitter. By measuring the resulting terahertz signal and employing a two-temperature model, we are able to calculate the dynamic electron temperature of the STE. This approach has an advantage over other methods, because it avoids the possibility of further heating of the sample by the probe pulse. This is especially significant for measurements of accumulative heat effects, which can dominate the optical damage mechanism when the spintronic terahertz emitter is pumped at a MHz repetition rate.
I Introduction
Designing efficient emission and detection devices in the terahertz regime ( to , between microwave electronics and infrared optics) has been a long-standing challenge. However, significant progress has been made in recent years, for example with the development of spintronic terahertz emitters (STEs) that offer much higher bandwidth than other commonly used emitters such as photoconductive antennas Kampfrath et al. (2013); Seifert et al. (2022); Torosyan et al. (2018); Ilyakov et al. (2022); Papaioannou et al. (2018); Nandi et al. (2019); Ilyakov et al. (2023); Seifert et al. (2016). STEs rely on the current-to-field conversion of a photocurrent originating from a femtosecond laser pulse excited spin current and spin-to-charge conversion in a magnetic heterostructure Seifert et al. (2022). They yield the advantage of a simple metallic thin-film-stack design which can be sputtered onto various substrates and therefore be fabricated and scaled at low cost Torosyan et al. (2018). Also, STEs can be pumped over a broad frequency range Ilyakov et al. (2022); Papaioannou et al. (2018); Nandi et al. (2019); Ilyakov et al. (2023), which enables a very adaptable implementation in table-top terahertz spectroscopy setups.
For sufficiently strong pump lasers, the limiting factor of the terahertz output power is the optical damage threshold. We have recently shown that the optical damage threshold of fiber-tip spintronic terahertz emitters can be classified into two regimes corresponding to two different destruction mechanisms Paries et al. (2024).
When an optical laser pulse hits the STE, it raises both the electron temperature in the metal layers and the lattice temperature. Within a few tens of picoseconds, the electron and phonon systems equilibrate to a new steady-state temperature, which is higher than the initial equilibrium temperature before the pulse. This steady-state temperature relaxes much more slowly – on the order of microseconds to milliseconds – due to thermalization with the environment trough coupling to the substrate Agarwal et al. (2021). As a result, when the time between successive laser pulses is short (e.g., at repetition rates above ), heat can accumulate in the system. This can lead to a progressive increase in the STE’s steady-state temperature over time. This gradual increase in steady-state temperature drives atomic interlayer diffusion, which is the dominant degradation mechanism of the STE at high repetition rates Paries et al. (2024).
To optimize thermal management in such cases, it is essential to understand the dynamics of both the electron and phonon systems in the STE. This includes, most importantly, the evolution of the electron temperature in the first few picoseconds following excitation, the electron–phonon coupling over the subsequent tens of picoseconds, and the longer-timescale heat transport from the pump spot to the surrounding area – both laterally within the STE layer and vertically into the underlying substrate. These thermal processes are investigated in the following using an optical-pump terahertz-probe setup.
The electron temperature in STEs consisting of a tri-layer structure of W(3 nm) | NiFe(3 nm) | Pt(3 nm) has previously been measured in a reflection geometry using optical-pump optical-probe spectroscopy for different pump fluences around . An increase in electron temperature above and relaxation dynamics on the order of several hundreds of ps back to a steady-state temperature were shown. Agarwal et al. (2021) However, in such measurements, it is unclear how much the electron temperature is influenced by the probe. Such effects are particularly confounding in the case of accumulated thermal effects, where the probed region does not return to equilibrium between consecutive pump pulses, as also mentioned by the authors Agarwal et al. (2021). Here we present a different approach for investigation of the electron temperature dynamics using an optical-pump terahertz-probe setup Beard, Turner, and Schmuttenmaer (2002) and a STE with a different ferromagnetic layer (). Due to the low photon energy of the probe pulse, it does not induce a significant change in the electron temperature of the STE, thereby eliminating a potential source of ambiguity in the measurement. Furthermore, the experiments are conducted at a low repetition rate of 1 kHz to enable an undisturbed observation of the electron temperature dynamics following the absorption of a single optical pump pulse. This approach ensures that cumulative heating effects from subsequent pulses, as discussed above, are avoided. While the electron–phonon dynamics observed here occur on picosecond timescales where optical probe-induced perturbations are typically small, the use of a terahertz probe inherently avoids any excitation via single- or multi-photon absorption. This provides a minimally invasive probing approach that may be especially important in future applications involving ultrafast dynamics, fragile materials, or resonant optical transitions. The estimated THz probe field strengths used here are approximately 100 kV/cm, which corresponds to a pulse energy of .
II Methods
[…]
The optical pump beam is blocked after the STE, and the two terahertz signals corresponding to the probe and the terahertz beam generated by the STE are detected with electro-optic sampling and balanced detection Koch et al. (2023), using the remaining of the laser power. Delay stages control the relative time delay between the arrival of the terahertz probe pulse and the STE pulse at the detector. Pairs of wire grid polarizers before and after the sample position are used to further control the polarization and intensity of the different beams.
The STEs are manufactured by depositing a W()()Pt() tri-layer structure onto thick sapphire substrates employing RF-diode sputtering. Details about the manufacturing process can be found in Ref. 15. Magnets are mounted around the STE and are used to control the polarization of the terahertz pulses emitted by the STE Seifert et al. (2022).
III Results and discussion
During measurements, the chopper is placed in the optical pump path. Because the terahertz probe beam is not modulated by a chopper, the corresponding signal is not amplified, and only the emission from the STE (induced by the pump pulse) is measured. Figure 2 shows typical measurement results, as well as a schematic drawing of the different relative time delays. In the top panel, the optical pump pulse reaches the STE wafer after the terahertz probe pulse; in the middle and bottom panel, the optical pump pulse reaches the STE wafer before the terahertz probe pulse with a different relative time delay between them. In the latter two cases, an additional peak is visible in the measured waveforms at the time delay of the terahertz probe pulse (highlighted in yellow). This indicates that the presence of the optical pump pulse modifies the signal strength of the terahertz probe pulse at the detector. To further investigate this effect, we rotate the polarization of the infrared detection beam with respect to the z-axis of the ZnTe detection crystal to suppress the STE signal contribution Planken et al. (2001). In this configuration, the chopper is still placed in the optical STE pump path, and the lock-in amplifier is referenced to that modulation. The measured signal therefore corresponds to optical STE pump-induced changes in the transmitted THz probe pulse.
Figure 3 shows a close-up view of the additional peak in the measured signal, along with a schematic representation of the different relative timing configurations of the optical pump pulses. When the optical pump pulse arrives after the terahertz probe pulse, shown in the top panel, no peak is visible. However, when the pump pulse excites the wafer before the terahertz probe pulse arrives (the second and all lower panels in Fig. 3), the additional peak appears in the time trace of the terahertz waveform, always at the same time delay as the terahertz probe pulse.
Following the panels from top to bottom, it is clear that the amplitude of the additional observable peak decreases with increasing time delay between the optical pump pulse and the terahertz probe pulse. This signal must originate from the influence of the optical pump pulse on the sample. The optical pump pulse, in addition to generating a THz signal, also produces an instantaneous increase of the electron temperature, which then relaxes on a picosecond time scale.Seifert et al. (2016); Wu et al. (2021); Vaitsi et al. (2024). The optical pump pulse therefore varies the transmittivity of the terahertz probe pulse through the STE. Since this change is modulated at the frequency of the optical chopper in the optical STE pump path, it is detected in the lock-in measurement of the THz waveform. Hayashi et al.Hayashi et al. (2022) demonstrated that changes in the optical reflectivity of metallic samples are sensitive to both electron and phonon temperatures, through modifications of the complex dielectric function. While an approximate proportionality between electron temperature and reflectivity can hold under specific conditions, the relationship is in general nontrivial and model-dependent.Hayashi et al. (2022) Since reflectivity and absorption both derive from the complex refractive index, they are intrinsically linked by the Kramers-Kronig relations. In our measurements, reflection is not negligible, and the observed change in transmission thus reflects a combined change in absorption and reflectivity. Thus, by measuring the change in transmission, we can track the variation in electron temperature induced by the optical pump. The amplitude of the observed additional peak as a function of the relative time delay between the optical pump pulse and the terahertz probe pulse is plotted in Figure 4a and 4b.
In Figure 4a, we observe a rapid initial drop, followed by a slower exponential decay with a time constant of . The exponential decay after the rapid initial drop is illustrated in the logarithmic plot in the inset of figure 4a where a linear behavior is visible. In the first (Figure 4b), we observe oscillations on a timescale of , superimposed on the slow decay. This corresponds to the transit time of the optical pump pulse through of sapphire (refractive index at ) Thomas et al. (1998). It is therefore likely that these oscillations originate from multiple reflections in the sapphire substrate, as illustrated in the inset in Figure 4b.
To understand the different timescales of the observed change in transmission, we consider the sequence of interactions occurring in magnetic materials after excitation by a femtosecond laser pulse. First, the energy of the optical pulse is absorbed by electrons, forcing the system out of equilibrium Bigot, Vomir, and Beaurepaire (2009). The electrons rapidly relax into hot thermalized populations on timescales below 100 fs Bigot, Vomir, and Beaurepaire (2009). These dynamics are faster than the duration of the probing terahertz pulse and are therefore not resolved in our measurements. However, they are responsible for the abrupt increase in amplitude seen in Figure 3b at pump–probe delays shorter than 1 ps. Subsequently, electron–phonon coupling becomes dominant. This interaction occurs over a few picoseconds and gives rise to the fast decay observed in Figure 3b within the first 5 ps after excitation. At this point, the electron and phonon systems have reached thermal equilibrium. The subsequent dynamics involve heat dissipation from the pump spot into the surrounding system.
This process involves two parallel channels: lateral heat transport within the STE layer, and vertical heat transfer across the STE–sapphire interface, which is limited by electron–interface scattering Agarwal et al. (2021); Guo et al. (2012). Although both processes occur simultaneously, they operate on different timescales. The in-plane heat diffusion is relatively fast on the timescale of a few tens of ps, while the interfacial heat transfer is slower due to thermal boundary resistance and on the timescale of a few hundreds of ps Agarwal et al. (2021). Together, these mechanisms contribute to the slower decay observed in Figure 3b, which extends to approximately 400 ps after excitation.
We note that our experimental time resolution does not allow us to capture the ultrafast rise of the electron temperature in the first tens of femtoseconds following optical excitation. As such, the peak electron temperature prior to electron–phonon equilibration is not directly accessible in our measurements. Our analysis is therefore limited to the evolution of the electron–lattice system after thermalization, and the estimated temperature values reflect this regime.
A quantitative description of these dynamics can be extracted from a simple two-temperature-model (TTM) as discussed in Ref. 10 and Ref. 23. The TTM describes the temperatures of the electronic system and the lattice, and , in a metal after an excitation of the electronic system. In our case, the excitation is the absorbed femtosecond pulse and the metal is the thin STE layer. In metallic thin films with thicknesses below the electron inelastic mean free path (IMFP), such as the ferromagnetic layers used in STEs, lateral transport of hot electrons is significantly suppressed. As a result, energy relaxation immediately following optical excitation is dominated by electron–phonon coupling rather than by hot-electron diffusion. Hohlfeld et al. (2000); Giri and Hopkins (2020) After excitation, the cooling of the electron system can be expressed with two coupled differential equations, which describe the heat conduction of the electrons and the lattice:Atxitia et al. (2010)
In these equations, and are the specific heats of the electrons and the lattice, respectively, is an electron-phonon coupling constant which connects the two equations and determines the rate of the energy exchange between the electrons and the lattice. is the ambient room temperature and is the heat diffusion time. The first term in equations (1) and (2) describes the coupling between the two systems. In equation (1), the second term represents the excitation source which is absorbed in the metal layer and excites the electron system. The last term describes the heat diffusion in the metal layer.
To solve these equations, we make several simplifying assumptions. Normally, the electron specific heat is strongly temperature dependent. Here, we use the simplification with , as proposed in Ref. 23. The temporal intensity of the excitation (pump) pulse is approximated as a Gaussian, and the absorbance of the laser pulse in the metal layer is approximated Koleják et al. (2024) as 32 % and assumed to be uniform over the illuminated area and depth in the metal layer. It is assumed that both the maximum lattice temperature and the equilibrium between electron and lattice temperatures are reached when the fast decay in Figure 4 ends and the slow decay starts, because the coupling between the systems vanishes as the temperatures approach each other Agarwal et al. (2021). We approximate the surface total specific heat of the STE layer as based on the thickness of the STE layer and an estimated specific heat capacity of the STE of based on the literature values for the specific heat capacity of platinum Yokokawa and Takahashi (1979) and tungsten Bronson, Chisholm, and Dockerty (1933). By taking into account the repetition rate of the laser (), and the average optical pump power (), as well as the pump spot area of , an absorbed fluence of is derived. The maximum temperature of the STE with the electron and the lattice temperature in equilibrium is then estimated to be . This maximum temperature enables calibration of the scale in Figure 5. It should be noted that the calculated absolute values have a large confidence interval, as the main parameters are approximated based on data of similar (but not identical) materials from literature.
The obtained fit, using the temperature-calibrated measurement data, is shown in Figure 5. The lattice temperature is displayed in green. The calculated electron temperature and the measured data are in very good agreement. However, the inset in Figure 5a shows that the oscillations due to the multiple reflections in the sapphire substrate are not captured by the model. From the fit to equations (1) and (2), a heat diffusion time of and an electron-phonon coupling constant of are obtained, which are directly connected to the electron-phonon scattering rate and are of the same order of magnitude as reported in Atxitia et al. (2010); Agarwal et al. (2021); Beaurepaire et al. (1996) for various comparable material systems.
Extending this model by modifying the excitation source to a sum of Gaussian pulses provides a solution to account for the multiple reflections in the sapphire substrate. The fit with this modified excitation source is shown in Figure 5b. Now, also the oscillations are in very good agreement with the fit. The small deviation of the measured data from the fit in the slow decay from to may be due to the excluded coupling term of the electronic system of the STE layer and the phononic system in the substrate Agarwal et al. (2021).
From our results, we find that the electron system relaxes to its initial state about after excitation. This suggests a threshold for the fluence-dominated damage process due to accumulated heat in the STE layer of around . In a previous study Paries et al. (2024), the optical damage process was connected to heat accumulation in the STE layer with a laser repetition-rate threshold of for the fluence of . This is two orders of magnitude lower than what we found here. The discrepancy can most likely be explained by the different material systems used in the study. The fiber-tip STE used in the previous study provided a much smaller surface area of the STE, leading to a saturated heat transfer to the surrounding area over time. Additionally, the heat transfer in the fiber-tip STE substrate (glass ) Simoncelli, Mauri, and Marzari (2023) is a factor of 30 weaker than the substrate used here (sapphire )Dobrovinskaya, Lytvynov, and Pishchik (2009), leading to a slower heat transfer away from the pump spot. This may explain the orders-of-magnitude higher repetition rate threshold in the current - very localized - experiment but can be investigated further. Recent studies have focused on optimizing STEs for higher terahertz output power. Limited heat transfer in the thin metallic STE layer was identified as one way to improve the optical damage threshold of the STE, which impacts the terahertz emissivity Vogel et al. (2022); Vaitsi et al. (2024).
IV Conclusion
In this work we presented a new method of measuring the electron temperature dynamics in STEs using a terahertz probe pulse. Our experiments show that optical-pump terahertz-probe measurements can be used to investigate the electron temperature dynamics and coupling mechanisms in STE structures. We find here that the electron system reaches a temperature of several 100 K after excitation. However, the maximum electron temperature can only be estimated due to the time resolution of the terahertz probe pulse. This method offers the distinct advantage that the electron temperature dynamics of the STE are not disturbed by the terahertz probe pulse, as it could easily be the case when using an optical probe pulse. The current method enables for example the study of the optical damage process described in Ref. 9, which is connected to heat accumulation in the STE layer. We find here that the electron system relaxes to its initial state after excitation after about . This suggests a threshold for the fluence-dominated damage process at around for tri-layer STEs on sapphire, much higher than previously found Paries et al. (2024). This difference is probably due to saturated heat transfer to the surrounding media. Further optical-pump terahertz-probe studies on the temperature relaxation could provide more quantitative insight on the thermodynamical processes causing heat accumulation and ultimately STE damage. A detailed investigation of the influence of substrate properties and pump spot size on the relaxation dynamics would offer further insight into the dominant cooling mechanisms in STEs. Such measurements, while beyond the scope of this work, represent a promising direction for future studies. Such studies may help to improve the thermal management of STEs, resulting in more efficient THz emitters. Furthermore, the presented method provides a minimally invasive probing approach that may be also interesting in future applications involving ultrafast dynamics, fragile materials, or resonant optical transitions.
Acknowledgements.
We wish to acknowledge funding from NSF ECCS-2300152, Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - TRR 173 - 268565370 (SPIN+X, project B11), and a Study Abroad Studentship of the German Academic Scholarship Foundation (Studienstiftung des deutschen Volkes).Authors Declarations
Conflict of Interest
The authors have no conflicts to disclose.
Author Contributions
Felix Selz: Data curation (lead); Formal analysis (lead); Investigation (lead); Methodology (lead); Writing-original draft (lead); Writing - review & editing (equal). Johanna Kölbel: Data curation (equal); Formal analysis (equal); Investigation (equal); Methodology (equal); Supervision (equal); Writing - review & editing (equal). Felix Paries: Methodology (equal); Writing - review & editing (equal). Georg v. Freymann: Conceptualization (equal); Funding acquisition (equal); Methodology (equal); Project administration (equal); Resources (equal); Writing – review & editing (equal). Daniel Molter: Conceptualization (equal); Funding acquisition (equal); Resources (equal); Writing - review & editing (equal). Daniel M. Mittleman Conceptualization (equal); Funding acquisition (equal); Methodology (equal); Project administration (equal); Resources (equal); Supervision (equal); Writing – review & editing (equal).
Data Availability Statement
The data presented in this study are available from the corresponding author upon reasonable request.